# Limitations of Existing Work, and Guidelines for Future Research

### v1

@@ 3.1 Dimensionality
Geographic locations are two-dimensional, considerably increasing complexity (McMillen, 2010). Furthermore, the choice of level of detail for spatial data is not straightforward. For example, a geographic area may be represented at the level of houses, blocks, neighborhoods, or census tracts. While level of detail may be application-dependent, a finer level of detail increases computational complexity (Haining, 2003). Dimensionality reduction can be tricky with big data, and fairness concerns add another layer of complexity by making it a multi-objective problem (i.e., reducing dimensionality while satisfying fairness constraints). Fair dimensionality reduction, a nascent area, is non-trivial, and an open problem. For example, Samadi et al. (2018) show PCA often has higher reconstruction error on one subgroup than another, e.g., women versus men, and this may occur even when the data contains a similar number of samples from both groups. They propose fair-PCA (Samadi et al., 2018), but it requires $d+1$ dimensions for 2 protected subgroups, and is prohibitively slow in practice, while Tantipongpipat et al. (2019) propose a polynomial-time Fair-PCA algorithm for $k=2$ subgroups. Spatial data’s correlations with many protected characteristics (e.g., race, ethnicity), with each characteristic having multiple subgroups, complicates fair dimensionality reduction in spatial decision-making. Techniques in other areas such as r-trees (Guttman, 1984) that are designed to take advantage of the unique relationships between the two spatial dimensions to improve performance for multi-dimensional spatial data can offer insights for leveraging unique spatial dynamics to improve upon generic dimensionality reduction methods for spatial scenarios.
@@ 3.3 Continuity of space
Geographic spaces are continuous in nature, and not as easy to analyze as discrete spaces (Xie et al., 2022). Discretizing spatial data is not straightforward due to features of spatial data such as the spatial structure, association, and heterogeneity (Cao et al., 2014; Haining, 2003). Furthermore, geographical partitionings are heavily context-dependent, and less likely to have partitionings that are universally well-accepted than other continuous variables such as income which has well-defined thresholds (e.g., for income tax). Additionally, when the decision-making scenario is spatio-temporal, i.e., involves travel time, there is one more continuous dimension – time, further complicating matters. Finally, achieving fairness with respect to continuous attributes is non-trivial. Most fair-AI work focuses on discrete variables (e.g., gender: male/female) (Mehrabi et al., 2021). Although recent work explores fairness with continuous attributes, many depend on computationally intractable statistical independence measures in practice (Jiang et al., 2022; Mary et al., 2019; Jha et al., 2021; Creager et al., 2019), lack estimation accuracy guarantees (Jiang et al., 2022; Cho et al., 2020; Roh et al., 2020), or are not sensitive to protected subgroups that are less likely to occur in the population (Jiang et al., 2022). The combined complexities of continuous in two- or three- (with time) dimensions spatial data and achieving fairness with respect to continuous attributes make this a non-trivial task.
3.4 Modifiable Areal Unit Problem
The modifiable areal unit problem (MAUP), a phenomena observed in geography, is a form of statistical bias that arises when point-based estimates of spatial phenomena are aggregated into spatial partitions; the results from such an analysis will be influenced by the shape and scale of the spatial partition (Wong, 2004; Fotheringham & Wong, 1991). For a given space, the manner in which we define geographic boundaries or vary the level of geographic partitionings may change the perception of fairness. In other words, data collected for the same region but at different spatial scales will give inconsistent or different results. A common example of MAUP in the real-world is gerrymandering, which we discuss in Section 5.
@@ 3.5 Spatial autocorrelation
– Waldo Tobler, the First Law of Geography Finally, there is the challenge of spatial autocorrelation. The First Law of Geography states that nearby things are more related than distant ones, violating the independence assumption in standard statistical tests, requiring careful selection of statistical methods. Recent works in ML study such spatial dependence in different fields. Deppner et al. (2022) propose a spatial cross-validation strategy to reduce spatial bias in tree-based algorithms commonly used for price and rent modeling in the real estate industry. In the geosciences, work explores spatial autocorrelation when spatial features like spatial lag are used in random forests (Liu et al., 2022), and incorporating spatial autocorrelation measures in ML modelling for predicting geotechnical information, such as ground surface elevation (Kim et al., 2023), and increasing the prediction accuracy of heavy metals in soil (Sergeev et al., 2019). However, none of these works look at spatial fairness or spatial correlation with respect to legally protected attributes. Recently, Majumdar et al. (2022) propose a hypothesis testing methodology to detect the presence of spatial autocorrelation and its strength in presence of correlation with a protected attribute, although they do not propose techniques to mitigate the observed spatial bias. Thus, despite significant practical applications, mitigating spatial autocorrelation and its interplay with protected attributes remains a deeply understudied area.
These complexities of spatial data hinder use of fair-AI work in spatial scenarios, motivating dedicated spatial fairness, as recent works have limitations hindering their adoption.
@@ 4 Limitations of current spatial fairness work
Initial attempts at spatial fairness definitions and techniques fall short in significant ways, as we detail below.
@@ 4.1 Spatial fairness definitions: Legal soundness
Many fairness definitions have been proposed in fair-AI research (Verma & Rubin, 2018; Mehrabi et al., 2021), but most are not directly applicable to spatial fairness due to the challenges specific to spatial data noted in Section 3. Moreover, many fairness definitions proposed so far will not hold up to legal scrutiny (Xiang, 2020; Wachter et al., 2021). They may not be deemed legal in a court of law, and may thus be inapplicable in practice. Worryingly, many definitions may be incompatible with U.S. anti-discrimination as well as E.U. non-discrimination law.
Incompatibility with anti-discrimination law may also be an issue for the handful of spatial fairness definitions proposed so far. For example, (Shaham et al., 2022) propose two types of spatial fairness: distance-based spatial fairness and zone-based spatial fairness. The former is concerned with decision-making scenarios based on an individual’s distance from a reference point, such as a shop running a marketing campaign offering discount vouchers. However, it is legal to make decisions based on an individual’s distance from a reference point. The latter definition, zone-based fairness, considers scenarios where an individual’s location is taken relative to specific zones. It is legal to make a decision based on an individual’s location relative to specific zones, unless the zone is directly correlated with a protected attribute, which current work does not distinguish. Indeed, school districts, banking and mortgage industries, et cetera use zones to determine school assignment or in the computation to determine mortgage rates. To ensure adoption in the real-world, it is imperative that future spatial fairness work is in compliance with legal definitions of anti-discrimination.
@@ 4.2 Spatial fairness techniques: limitations
4.2.1 Do not close the loop
Perhaps the most concerning limitation of most spatial fairness techniques so far is that current research on spatial fairness does not yet close the loop: They do not demonstrate that their proposed spatial fairness methods addressed unfairness present due to location’s correlation with protected characteristics (Shaham et al., 2022; Sacharidis et al., 2023). Although discrimination on the basis of legally protected characteristics is the motivation for the spatial fairness techniques proposed so far, there are no experiments demonstrating techniques that can reduce discrimination on the basis of such attributes, like race.
The incompatibility of most spatial fairness definitions with U.S. and E.U. laws may hinder their practical adoption. Moreover, these techniques must unequivocally demonstrate improvement in fairness with respect to protected characteristics before being considered for adoption.
@@ 4.2.2 Take agency away from people
Some of the formulations of spatial fairness proposed so far use a fuzzy boundary instead of hard boundaries as are typically employed in domains such as education (e.g., school districts), banking, or housing. While this may have some technical advantage, a significant consequence of this would be to take agency away from people. For example, with clear school-district boundaries, a low-income family with school-age children can choose to make sacrifices to live in a better school district (Hristov & Khare, 2016). If such boundaries are fuzzy, however, then this can impact the decision-making ability of individuals and families in a deeply negative manner. Wealthy families will still be able to ensure a good education for their children by living squarely in the heart of good school districts or simply paying for private schools. But lower and middle income families will suffer due to the opaqueness of the fuzzy boundary, and not being able to assess the chance that a particular address may have of securing their children admission in a good school. Research has shown that one of the barriers for disadvantaged families to be able to benefit from public benefits they are eligible for is lack of knowledge of eligibility rules (Wu & Eamon, 2010; Anderson, 2002; Stuber & Kronebusch, 2004; Zedlewski, 2003) and complicated application processes (Wu & Eamon, 2010; Morrow & Horner, 2006; Zedlewski et al., 2002). Therefore, increasing the complexity of assessing which school district an address may belong to will overwhelmingly harm those least likely to be equipped to deal with it efficiently.
@@ 4.2.3 Do not build upon existing knowledge
Research is about building on existing work to stand on the shoulders of giants. A considerable amount of work explores different aspects of spatial fairness with respect to protected attributes. For example, literature in spatial scan statistics (Wong, 2004; Fotheringham & Wong, 1991) and transport geography (Guo & Bhat, 2004; Mitra & Buliung, 2012) explore MAUP from different angles, which can offer ideas for how to grapple with MAUP in spatial fairness. Similarly, work in public policy studies spatial segregation patterns with respect to race, ethnicity, and immigration in housing and schools across the country (Schill & Wachter, 1994; Sanchez, 2006; Saiz & Wachter, 2011; Schill & Wachter, 1995). Research in economics can help offer guidance on which spatial measures help improve economic development, especially for the most disadvantaged (Chetty et al., 2016; Brata et al., 2009; Glasmeier, 2018)
@@ 5 Guidelines and Future Directions
We define guidelines to help future spatial fairness research avoid the pitfalls and challenges outlined in Sections 3 and 4, and delineate future directions.
5.1 Close the loop
Research exploring spatial fairness so far does not close the loop, i.e., demonstrate that their proposed work reduces negative effects on people belonging to protected classes. Spatial fairness work should depict empirical experimental evaluation of the extent of spatial bias with respect to protected attributes in the problem domain with and without the proposed methodology and its impact on other problem objectives (e.g., prediction accuracy; efficiency of matching students to schools). For example, suppose a methodology places fairness constraints on location $L(x)$ to mitigate its bias associated with protected attribute $x$, getting $L^{\prime}(x)$ in the process. Experiments must demonstrate the extent of bias between $L(x)$ and $L^{\prime}(x)$ with $x$, and the trade-offs (if any) of mitigating bias with accuracy on outcomes, in addition to comparisons with previously proposed methods.
5.2 Avoid disparate impact
Under U.S. laws, practices that unintentionally harm people of a protected characteristic, such as in employment and housing, may be considered discrimination (Corbett-Davies & Goel, 2018; Feldman et al., 2015), or disparate impact in legal terms. Research has shown the potential for computational decision-making models to perpetuate disparate impact (Barocas & Selbst, 2016). To address this, a growing body of work advocates for the development of disparate impact assessments, potentially modeled on similar analyses in industries that must comply with non-discrimination laws (MacCarthy, 2017). Other work also argues for such vigilance or assessment to continue after model deployment, perhaps by external auditors, to ensure models function as intended and have no unpleasant unintended consequences (Reisman et al., 2018; Saxena et al., 2023a). Assessing disparate impact in computational systems is not trivial, with healthy debate in the fair-AI community about how to detect and assess disparate impact. Work has proposed tests and methods to detect and measure disparate impact in data-driven systems (Feldman et al., 2015; Mehrabi et al., 2021; Metcalf et al., 2021; Reisman et al., 2018), and critiqued tests and their misalignment with U.S. anti-discrimination law (Selbst et al., 2019; Xiang & Raji, 2019). Assessing disparate impact in spatial decision-making, however, must incorporate the unique challenges of location and spatial assessments of outcomes with respect to protected groups.
In domains like employment, credit, and housing, tools that measure such disparate impact and are compliant with U.S. and E.U. anti-discrimination laws would likely enjoy great demand by supporting companies in prevention, and empowering individuals to identify if they have been subject to unfair effects of discriminatory practices.
5.3 Location: a protected or immutable characteristic
Immutability is deeply associated with anti-discrimination law in the U.S. (Hoffman, 2010). Recent interpretations are a characteristic “that either is beyond the power of an individual to change,” [In re Acosta, 19 I. & N. Dec. 211, 233-34 (B.I.A. 1985)] or in other terms, “either unchangeable in absolute terms or so fundamental to identity or conscience that individuals effectively cannot and should not be required to change it” [In re Acosta, 19 I. & N. Dec. 211, 233-34 (B.I.A. 1985).] or “changing it would require great physical difficulty, such as requiring a major physical change or a traumatic change of identity.” [Watkins v. U.S. Army, 875 F.2d 699, 726 (9th Cir. 1989)] (Hoffman, 2010), thereby establishing an immutable characteristic need not be unchangeable to be considered immutable under anti-discrimination law. This, in addition to ongoing efforts to afford protection against discrimination on the basis of zip codes, brings us to the next guideline: consider location a protected attribute, or at least an immutable one. We outline example use cases below.
In domains with demonstrated history of significant location-related discrimination, zip codes could be considered a joint protected attribute with the legally protected attribute it perpetuates bias against. For example, an intuitive measure of individual fairness is to treat similar people similarly (Dwork et al., 2012). With location and race as a joint protected attribute, two applicants who are similar to each other on non-protected aspects should receive similar decisions, regardless of their race and location. When impractical, location should be treated as an immutable attribute. For example, some fair- and explainable-AI research aims to provide applicants with actionable points they can improve to increase their likelihood of a positive decision (Ustun et al., 2019). Suggested actions should, however, be reasonably achievable. For example, one can recommend increasing annual salary by \$2,000, but if the salary of an applicant with a PhD degree is low for someone with a PhD degree, one cannot recommend them to undo their PhD. In a similar vein, for low-income individuals and families, residence should be treated as immutable, i.e., something an applicant of a particular socio-economic status cannot reasonably change since they commonly do not have the freedom to make this choice, sometimes even if they can afford to (e.g., due to source of income discrimination (see Sec 2.1)).
5.4 Be resistant to MAUP
MAUP is a dilemma unique to spatial analyses (Sec 3.4). At a high level, two types of spatial biases can distort results when analyzing spatial data. The scale effect occurs when point data is aggregated at different levels of aggregation. Despite using the same point data, different aggregation levels can represent different patterns. On the other hand, the zonal effect occurs when spatial is grouped by different artificial boundaries, leading to different results. Research over the years has found MAUP to exhibit a high level of unpredictability in terms of its intensity and effects in multivariate statistical analyses, significantly surpassing the challenges associated with MAUP in univariate or bivariate statistical analysis (Fotheringham & Wong, 1991)
A classic real-world example that can highlight the dangers of MAUP is the issue of gerrymandering (Nelson & Brewer, 2017; Manley, 2021), the practice of redrawing political boundaries or election districts to serve the interests of a particular political party. There is work in the fields of spatial statistics and geography for methods to address MAUP generally (Tuson et al., 2019; Hennerdal & Nielsen, 2017) or in specific applications such as traffic safety (Xu et al., 2018) or designing zonal districts without falling into the MAUP trap that research on spatial fairness can potentially draw from (Martin, 2003; Tuson, 2022).
5.5 Different definitions for public and private sectors

### v2

@@ 3.1 Dimensionality
A critical distinction between spatial fairness and typical fair-AI work lies in the cardinality of the protected attribute. Fair-AI research considers attributes like race or gender, which possess a limited number of distinct values. However, when location is treated as a protected attribute, the number of values can expand to thousands (e.g., zip codes).
This high cardinality presents a fundamental barrier for existing dimensionality reduction techniques. For instance, Samadi et al. (2018) propose Fair-PCA, but their analysis shows that if a protected attribute has $k$ possible values, the algorithm requires a target subspace of $k+1$ dimensions to satisfy fairness constraints. While this is feasible for binary attributes ($k=2$), it is intractable for spatial data. With thousands of location values, the required dimensionality would effectively negate the purpose of reduction. Consequently, generic fair dimensionality reduction methods are ill-suited for spatial scenarios where the protected attribute is defined by fine-grained geographic units.
Fair dimensionality reduction is non-trivial and spatial data’s correlations with many protected characteristics (e.g., race, ethnicity, national origin), with each characteristic having multiple subgroups, complicates fair dimensionality reduction in spatial decision-making even more. Techniques in other areas such as r-trees (Guttman, 1984) that are designed to take advantage of the unique relationships between the two spatial dimensions to improve performance for multi-dimensional spatial data can offer insights for leveraging unique spatial dynamics to improve upon generic dimensionality reduction methods for spatial scenarios.
@@ 3.3 Continuity of space
Geographic spaces are continuous in nature, and not as easy to analyze as discrete spaces (Xie et al., 2022). Discretizing spatial data is not straightforward due to features of spatial data such as the spatial structure, association, and heterogeneity (Cao et al., 2014; Haining, 2003). Moreover, geographical partitionings are heavily context-dependent, and less likely to have partitionings that are universally well-accepted than other continuous variables such as income which has well-defined thresholds (e.g., for income tax). Additionally, when the decision-making scenario is spatio-temporal, e.g., involves travel time, there is one more continuous dimension – time, further complicating matters. Finally, achieving fairness with respect to continuous attributes is non-trivial. Most fair-AI work focuses on discrete variables (e.g., gender, race) (Mehrabi et al., 2021). Although recent work explores fairness with continuous attributes, many depend on computationally intractable statistical independence measures in practice (Jiang et al., 2022; Mary et al., 2019; Jha et al., 2021; Creager et al., 2019), lack estimation accuracy guarantees (Jiang et al., 2022; Cho et al., 2020; Roh et al., 2020), or are not sensitive to protected subgroups less likely to occur in the population (Jiang et al., 2022). The combined challenges of continuous spatial data in two or three dimensions (with time) and fairness with continuous attributes make it a nontrivial task.
3.4 Modifiable Areal Unit Problem (MAUP)
Identified in 1934 (Gehlke and Biehl, 1934) and coined as MAUP in 1979 (Openshaw, 1979), it is a type of statistical bias observed in geography. At a high level, two types of biases can distort results when analyzing spatial data. The scale effect occurs when point data is aggregated at different levels of aggregation. Despite using the same point data, different aggregation levels can represent different patterns. In contrast, the zonal effect occurs when spatial is grouped by different artificial boundaries, leading to different results. In other words, data collected for the same region but at different spatial scales will give inconsistent results. Research has found MAUP can be highly unpredictable in multivariate statistical analyses, with effects that are often substantially more severe than those observed in univariate or bivariate analyses(Fotheringham and Wong, 1991). A common real-world example of MAUP is gerrymandering.
@@ 3.5 Spatial autocorrelation
– Waldo Tobler, the First Law of Geography
Finally, there is the challenge of spatial autocorrelation: nearby things are more related than distant ones. This violates the independence assumption in many standard statistical tests, requiring careful selection of statistical methods. Recent works in ML study such spatial dependence in different fields. Deppner et al. (2022) propose a spatial cross-validation strategy to reduce spatial bias in tree-based algorithms commonly used for price and rent modeling in the real estate industry. In the geosciences, work explores spatial autocorrelation when spatial features like spatial lag are used in random forests (Liu et al., 2022), and incorporating spatial autocorrelation measures in ML modelling to predict geotechnical information like ground surface elevation (Kim et al., 2023), or increasing the prediction accuracy of heavy metals in soil (Sergeev et al., 2019). However, none of them look at spatial correlation with respect to legally protected attributes. Majumdar et al. (2022) propose a hypothesis testing framework to detect spatial autocorrelation and its strength in presence of correlation with a protected attribute, although they do not propose techniques to mitigate the observed bias. Thus, despite significant practical applications, mitigating spatial autocorrelation and its interplay with protected attributes remains deeply understudied.
These complexities hinder use of fair-AI work in spatial settings, motivating dedicated spatial fairness research.
@@ 4 Limitations of current spatial fairness work
However, early work on spatial fairness definitions and techniques fall short in significant ways, as we detail below.
@@ 4.1 Spatial fairness definitions: Legal soundness
Many fair-AI fairness definitions will not withstand legal scrutiny (Xiang, 2020; Wachter et al., 2021), making them inapplicable in practice. Worryingly, they may be incompatible with U.S. anti-discrimination and E.U. non-discrimination law. Incompatibility with anti-discrimination law may also be an issue for the handful of spatial fairness definitions proposed so far. Shaham et al. 2022 propose two definitions: distance-based and zone-based spatial fairness. The former focuses on decision-making scenarios based on how far an individual is from a reference point, such as a shop offering discount vouchers to customers within a certain distance. The latter considers decisions based on whether an individual falls within particular geographic zones. However, both definitions only consider location while evaluating fairness, without accounting for legally protected attributes. This substantially limits their real-world applicability, since making decisions based on location alone is not unlawful. Related work (Sacharidis et al., 2023; Xie et al., 2022) similarly considers fairness only with respect to location, limiting their practical usefulness. To ensure adoption in the real-world, it is imperative that future spatial fairness work is aligned with established legal frameworks for anti-discrimination.
@@ 4.2 Spatial fairness techniques: limitations
4.2.1 Do not demonstrate reduction in bias
A critical limitation of current spatial fairness techniques is the failure to “close the loop.” While these methods are motivated by location’s correlation with protected characteristics, they rarely validate that their spatial interventions actually achieve this goal (Shaham et al., 2022; Sacharidis et al., 2023): they lack experiments showing that addressing location-based bias successfully mitigates unfairness toward the underlying protected groups. To be considered for adoption, techniques must unequivocally demonstrate they result in tangible improvements for the protected characteristics.
@@ 4.2.2 Take agency away from people
Some formulations of spatial fairness proposed so far use a fuzzy boundary instead of hard boundaries as are typically employed in domains such as education (e.g., school districts), banking, or housing. While this may have some technical advantage, a significant consequence of this would be to take agency away from people. For example, with clear school-district boundaries, a low-income family with school-age children can choose to make sacrifices to live in a better school district (Hristov and Khare, 2016). If such boundaries are fuzzy, however, then this may impact the decision-making ability of families in a deeply negative manner. Wealthy families will still be able to ensure a good education for their children by living squarely in the heart of good school districts or simply paying for private schools. But lower and middle income families will suffer due to the opaqueness of the fuzzy boundary, and not being able to assess the chance that a particular address may have of securing their children admission in a good school. Past research shows that a big barrier to disadvantaged families benefiting from public benefits they are eligible for is lack of knowledge of eligibility rules (Wu and Eamon, 2010; Anderson, 2002; Stuber and Kronebusch, 2004; Zedlewski, 2003) and complicated application processes (Wu and Eamon, 2010; Morrow and Horner, 2006; Zedlewski and others, 2002). Therefore, fuzzy boundaries may disproportionately harm those who are least equipped to navigate the process efficiently.
@@ 4.2.3 Do not build upon existing knowledge
Research is about building on existing work to stand on the shoulders of giants. Considerable work explores various aspects of spatial fairness. For example, literature in spatial scan statistics (Wong, 2004; Fotheringham and Wong, 1991) and transport geography (Guo and Bhat, 2004; Mitra and Buliung, 2012) explore MAUP from different angles, which can offer ideas for how to grapple with MAUP in spatial fairness. Similarly, work in public policy studies spatial segregation patterns with respect to race, ethnicity, and immigration in housing and schools across the country (Schill and Wachter, 1994; Sanchez, 2006; Saiz and Wachter, 2011; Schill and Wachter, 1995). Economic research can offer guidance on which spatial measures help improve economic development for the disadvantaged (Chetty et al., 2016; Brata and others, 2009; Glasmeier, 2018).
@@ 5 Guidelines and Future Directions
To mitigate spatial bias, we propose a holistic rethinking of the modeling pipeline and establish guidelines from foundational steps to post-deployment monitoring
5.1 Foundational Steps
5.1.1 Establish socio-technical context
Clearly define domain-specific constraints and requirements. Before technical development, researchers must document the legal, regulatory, and public policy constraints governing decision-making systems in that domain. In other words, document the constraints or requirements the system is expected to satisfy. In addition, identify the population(s) likely to be impacted by the decision-making system, and the subgroups most vulnerable to marginalization.
5.1.2 Data coverage and representativeness
Evaluate training data for gaps in coverage and demographic representativeness. Location data collection is often systematically biased, yielding samples that reflect only a subset of the underlying population due to skewed data sources and collection practices. For instance, location data collected from mobile applications may underrepresent certain demographics, as device type (e.g., iPhone vs android) and app usage patterns vary significantly by socioeconomic status and age (Zeighami and Shahabi, 2024). Researchers should explicitly document these gaps and any reweighting or other procedures used to mitigate them.
5.1.3 Engage legal, community stakeholders
Collaborate with experts and affected communities. Identify stakeholders, including experts who shape, interpret, or enforce the legal, regulatory, and public policy constraints in the decision-making domain. They can aid computer scientists translate the normative requirements into precise technical specifications and evaluation criteria, enabling verification that system outputs meaningfully satisfy the constraints. Moreover, in high-impact domains like housing, community stakeholders, especially those underrepresented in the data, should be consulted be consulted during development and evaluation. This sustained engagement can uncover factors and values important to the general public that may not be captured in formal requirements, like distinct school district boundaries so communities retain meaningful agency in how decisions affect them.
5.2 Model Design Principles for Spatial Fairness
5.2.1 Tailoring Fairness to Domain and the Law
Develop domain-specific fairness definitions. Spatial fairness definitions must be tailored to the specific problem domain in collaboration with stakeholders identified in the foundational phase. Legal experts are particularly vital to ensure the proposed definition: (1) strictly complies with the domain-specific anti-discrimination laws; (2) prevents the inadvertent encoding of discriminatory intent (i.e., disparate treatment); and (3) enables evaluations to meaningfully assess whether system outputs avoid disparate impact.